OpenStudy (anonymous):

The price of movie tickets has increased over time with inflation. Let's let the equation y = 0.50(1.06)x represent the price of movie tickets for the years after 1950. Using this equation, fill in the following T-table for the years 1950, 1960, 1970, 1980, and 1990. Use X = 0, 10, 20, 30, and 40 to represent these years (the number of years after 1950). Then type in your answers in the space given. A table where the top row or x represents number of years after 1950 and the bottom row or y represents Ticket Price in dollars. X = 0 10 20 30 40

OpenStudy (anonymous):

If we make a function \[f(x)=0.5(1.06)^{x}\] and then solving for each of the values we'll get: 1950 - f(0)=0.5 $ 1960 - f(10)= 0.9 $ 1970 - f(20)= 1.6 $ 1980 - f(30)= 2.9 $ 1990 - f(40)= 5.4 $ You can look at the graph here: http://www.wolframalpha.com/input/?i=0.5%281.06%29^x+x+from+0+to+40

OpenStudy (anonymous):

To clarify, when i write 1950 - f(0) the "-" does not represent a minus.

OpenStudy (anonymous):

I FOUND YOU! ive been looking for you :)